OpenStudy (anonymous):

(97)3 into a single exponent remember 7 and 3 are exponents

OpenStudy (anonymous):

\[(9^7)^3=9^{21}\]

OpenStudy (anonymous):

hickninja is incorrect

OpenStudy (stormfire1):

(9^7)^3 = 9^(7*3) = 9^(21)

OpenStudy (stormfire1):

hick is correct

OpenStudy (anonymous):

A 9(7+3) b 9(7x3) c 97x93 d97+93

OpenStudy (stormfire1):

9^(7*3)

OpenStudy (anonymous):

\[(9^7)^3 = 9^7 \times 9^7 \times 9^7\]\[=9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9 \times 9\]\[=9^{21}\]

OpenStudy (anonymous):

thank u guys very much

OpenStudy (stormfire1):

The rule is, when you raise an exponent to another exponent, you multiple the exponents and keep the same base.